there are 400 students enrolled in euclid middleschool .of these 85 study spanish and 50 study french.if 120 students study either spanish or french .how many students study both french and spanish?
Answer by ikleyn(52879)

.

The answer is  85 + 50 - 120 = 135 - 120 = 15.



Explanation


I do not need to explain why I added 85 and 50 - it is clear.


I need only to explain why I subtracted 120.


    It is because when I added 85 and 50, I exceeded the value of 120.


    I exceeded it exactly by the value of those who study both languages.


    Therefore, when I subtracted 120, I exactly got the number of those who study both languages.


Answered, solved, explained and completed.

---------------

See the lesson
    - Counting elements in sub-sets of a given finite set
in this site.